# Revealing Nanostructures in High-Entropy Alloys via Machine-Learning Accelerated Scalable Monte Carlo Simulation

### v1

@@ 2.2 Energy model
Figure 2: The accuracy of the energy model and the calculated specific heats for FeCoNiAlTi and MoNbTaW. (a) The training error of the EPI model for FeCoNiAlTi. (b) The validation error of the EPI model for FeCoNiAlTi (0.7/0.3 train-valid splitting). (c) The training error of the loca_SRO model for FeCoNiAlTi. (d) The validation error of the local_SRO model for FeCoNiAlTi. (e) The specific heats $C_{v}$ of FeCoNiAlTi calculated with different supercell sizes. (f) The specific heats of MoNbTaW calculated with a 1 million atom supercell.
@@
In this work, we focus on two types of energy model: effective-pair-interaction (EPI) model [60, 48], and a nonlinear model that uses the local SRO parameter as the input feature. The EPI model is a generalized Ising model that constructed via machine learning from the DFT data. By automatically selecting the interaction range via Bayesian information criterion (BIC) [48], the EPI model demonstrates that it can predict the DFT-calculated configuration energies of a series of HEAs with very high accuracy and has been applied to study the order-disorder transitions in HEAs [33]. The local-SRO model, on the other hand, generalizes the EPI model by adding a quadratic term. As will be shown in the following discussion, this simple nonlinear term has a profound impact on the computing pattern, and serves as a prototype for more involved machine learning models. A detailed description of the two energy models are present in the Method 3.2.
@@
Using a train-validation splitting of 70% and 30%, we evaluated the accuracy of the second-nearest-neighbor (2nd-NN) EPI model and the loca_SRO model and the results are shown in Fig. 2 (a-d). Both models are trained with the Adam optimizer and back-propagation algorithm, as will be detailed in the method section. It can be seen that both of the two models demonstrates very high accuracy. For the EPI model, the validation root-mean-square-error (RMSE) is 0.1819 mRy, or approximately 2.5 meV, which is well within quantum chemical accuracy (approximately 3.16 mRy/atom), as well as smaller than the typical errors of $5-12$ meV/atom in the DFT methods [61]. The local_SRO model demonstrates slightly higher accuracy than the EPI model, with a validation RMSE of 0.1643 mRy/atom, which can be attribute to its inclusion of high-order interactions. The high accuracy of the models is also demonstrated in the $R^{2}$ scores, which are all higher than 0.995. The errors from the train and validation datasets are also very close to each other, indicating the absence of overfitting in the models. For comparison, we notice that this accuracy is higher than that in Ref. [29], which reported as 14.7 meV/atom by applying a similar pair-interaction model in the FeCoNiAlTiB system. We believe one reason for the enhanced accuracy is our approach to obtain the DFT dataset, which combines both random configurations, as well as configurations from Monte Carlo simulations, as described in Ref. [33].
@@
Using the trained models, we calculated the specific heats $C_{v}$ in the FeCoNiAlTi HEA in temperatures range from 50 K to 2000 K. The results are shown in Fig. 2 (e). It can be seen that there is a sharp order-disorder transition at a low temperature of about 300 K. The sharp peak evolves to singularity with the increase of the system size, which signifies the occurrence of a first-order transition. We also calculated the $C_{v}$ curve for the MoNbTaW refractory HEA. The results in Fig. 2 (f) agree well with the results in Ref. [33], which are calculated using a small system of 1000 atoms. While the order-disorder transition at lower temperatures (between 250 K and 500 K) in the $C_{V}$ curve cannot be directly compared with experiments due to the ignorance of kinetic barrier effects in the lattice MC method, the results at elevated temperatures are more suitable for comparison if experimental data are available, as demonstrated in Ref. [34]. A comparison of Fig. 2 (e) and (f) in the temperature range between 500 K and 2000K shows an interesting difference between the two HEAs: The $C_{v}$ curve in $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ is generally flat, but demonstrates clear fluctuation as the temperature changes. By comparison, the $C_{v}$ curve of MoNbTaW in the same temperature region is smooth, but with an obvious order-disorder transition around 1000 K (the exact location depends on the energy model and different values have been reported in various theoretical works [62, 28, 33, 63]). We think that such a difference originates from their distinct chemical concentrations: $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ is nonequimolar while MoNbTaW is equimolar. In equimolar systems, the ordered phase often has a well-defined stoichiometry (e.g., AB in a binary alloy). The transition between ordered and disordered phases is typically sharper and more likely to be first-order. In non-equimolar systems, the ordered phase may deviate from ideal stoichiometry, leading to defects or partial ordering. We will present more results in the following discussion of the nanostructures.
2.3 Nanostructures in FeCoNiAlTi
@@
Figure 3: Simulation results for FeCoNiAlTi with 1 million atoms. (a) A snapshot of the the (001) face at 1000 K. (b) a perspective 3D snapshot of the configuration at 1000 K, from the signified direction (c) A snapshot view of the (001) face at 2000 K (d) a 3D view of the configuration at 2000 K. (e) High-resolution atom maps showing the atomistic distribution within the $L1_{2}$ nanoparticles of the Al7Ti7 alloy. Reproduced from Fig. 2 in Ref. [39]. (f) TEM image of the Al7Ti7 alloy showing the nanostructured morphology. Reproduced from Fig. 1 in Ref. [39] with reprint permission from Science.
@@
Using the local SRO model, we first simulated the chemical phase changes of the $\mathrm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ fcc alloy using a supercell of approximately one million atoms. This alloy has demonstrated exceptional combined strength and ductility in experiment, and this attractive property has been attributed to the formation of nanoparticles in the FeCoNi matrix [39]. In experiment, the alloy sample is homogenized at 1150 ${}^{\circ}C$ (1423 K)for 2 hours and aged at 780 ${}^{\circ}C$ (1053 K) for 4 hours. In the canonical MC simulation, the lattice dimension is chosen as $72\times 60\times 60$, therefore, the total number of atoms is $72\times 60\times 60\times 4=1,036,800$. We initialize the configurations randomly at 2000 K, then decrease the simulation temperature by 50 K each time till 1000 K, which is approximately the temperature that the samples being aged. At each temperature, we skip $4\times 10^{5}-1.5\times 10^{6}$ MC sweeps before recording the following 20,000 sweeps for calculating the specific heats $C_{v}$ in Fig. 2 (e). The configurations at 2000 K and 1000 K are demonstrated in Fig. 3. It can be seen that the system is close to random at 2000 K. As the temperature decreases to 1000 K, the nanoparticles stabilize. The distribution of elements in the nanoparticle, as shown in Fig. 3 (a-b), reveals that it forms a slightly randomized A3B $L1_{2}$ structure, with Ni and Co occupying the A sites primarily, Al and Ti occupying the B sites and Fe can occupy either A or B sites. This agrees well with the experimental results in Ref. [39], as reproduced in Fig. 3 (e-f). The diameter of the cluster shown in Fig. 3 (a) is about 17.8 nm, which is in agreement with the experimental observation shown in Fig. 3 (f). Similar results are also reported in Ref. [64] for the $\rm{Fe_{25}Co_{25}Ni_{25}Al_{15}Ti_{10}}$ HEA, where the fcc phase consisted of a $\gamma$ Fe-(Co,Ni)-based solid-solution matrix (A1), and coherent primary $\gamma^{\prime}$ (Ni,Co)3-(Ti,Al)-based intermetallic $L1_{2}$ precipitates.
@@
Figure 4: Comparing the simulated configurations of the 1 billion atom FeCoNiAlTi with the experimental results. (a) A histogram for all clusters of sizes larer than 100 atoms. Five clusters of different sizes are selected and shown as vertical dashed lines. (b) The chemical concentrations of the matrix (cluster_0) and cluster_1, as well as the rest clusters. (c-f) Snapshots of the [001] face of NiCoFeAlTi at T=1000K, with (c) as the enlarged upper-right corner of (d), and (e) as the enlarged upper-right corner of (f). (c-d) show the different elements, and (e-f) show the atomic local energies. (g-i) 3D snapshot of cluster_4, 3, and 2, respectively.
From Fig. 3, it is obvious that even the 1 million-atom supercell is not large enough for studying the nanostructures in the FeCoNiAlTi HEA. To investigate the microstructures of the FeCoNiAlTi HEA at larger length scale, we employs a $630\times 630\times 630$ fcc supercell, in which the total number of atoms is 1,000,188,000. Due to the huge supercell size, we use the faster EPI model rather than the local SRO model. Similar to the 1M atoms case, we evenly decrease the simulation temperature from 2000 K to 1000 K using a step of 50 K. For the 1B atom system, we no longer record the MC configurations and run $10^{5}$ MC sweeps before recording the configuration at $T=1000$ K. Note that, generally speaking, the $10^{5}$ number of MC sweeps is not large enough for a supercell as large as one billion atoms to reach thermal equilibrium, so the simulation is closer to simulated annealing. Nevertheless, this nonequalibrium does not necessarily present as a problem since the actual synthesis of alloys is typically a nonequilibrium process involving rapid heating-cooling process, in which the phases at high-temperature can be trapped by kinetic barrier and be maintained to low temperature.
@@
The simulation results for this one billion-atom system are presented in Fig. 4. To study the size, composition, and morphologies of the nanoparticles, we also employ the union-find algorithm to identify the nanoparticles, as will be detailed in the method sections. We refer to all sets of atoms identified by the union-find algorithm as clusters, which can be of sizes ranging from a single isolated atom to the whole matrix phase. We make a histogram of all clusters of sizes larger than 100 atoms in Fig. 4 (a). It can be seen that the largest one is cluster_0, which contains about 0.6 billion atoms. This is the matrix phase, which contains only Fe, Co, and Ni, as demonstrated in Fig. 4 (b). Note that the absence of Al and Ti atoms in cluster_0 is a result of the union-find algorithm, which automatically identifies isolated Al and Ti atoms in cluster_0 as individual clusters of size one. From Fig. 4 (c), we see that the FeCoNi matrix phase is a disordered A1 structure. The cause of the disorder can be seen from Fig. 4 (e) and (f), which show the local energies of each atom with the color bar. It can be seen that in the matrix phase, the local energies of different atoms are generally close to each other, which means that the Fe, Co, and Ni atoms have no significant site preferences. On the other hand, the nanoparticles, which forms $L1_{2}$ structure as shown in Fig. 4 (c), contains sites with much lower atomic local energies, as shown in Fig. 4 (e). In fact, the local energy difference is exactly the criteria we make use of to distinguish the matrix and nanoparticles, as described in Method 3.4. The FeCoNi matrix phase is the entropy-stablized disordered phase, while the $L1_{2}$ phase is the enthalpy favorable ordered phase. In other words, the occurrence of both disordered A1 structure and ordered $L1_{2}$ structure is a result of the competition between the entropy and enthalpy contributions in the free energy.
@@
To study the size, compositions, and morphologies of the nanoparticle in more detail, we choose four representative nanoparticles, as illustrated in Fig. 4 (a) as cluster_1, 2, 3, and 4, which have sizes of 264,792,042, 1,007,916, 102,389, and 20,937 atoms. Their chemical concentrations are also listed in Tab. 2, which aligns excellently with the experimental results [29]. From Fig. 4 (a), we note that the peak of the histogram is at about $10^{5}$ atoms, which corresponds to a cluster of 29 atoms, or 10.3 nm in diameter. For cluster_1, similar calculation gives a value of 22.0 nm in diameter, if we assume that the particles are spherical. However, inspecting the 3D morphology of the clusters reveals that they are actually not all simple spheres, as shown in Fig. 4 (g-i): While sphere is a good approximation for the smaller cluster_4, the larger cluster_3 and cluster_2 are both comprising of multiple nanoparticles, which indicates that the smaller nanoparticels can merge together to form larger ones, as shown by the dendritic structures at the surface of the nanoparticles. In fact, from Fig. 4 (a,b) we see that the largest $L1_{2}$ cluster: cluster_1, contains a total of more than 264 million atoms, which accounts for 66% of the atoms in the nonmatrix phases. In other words, most of the nanoparticles shown in Fig. 4 (c-f) are actually connected at 3-dimension and are parts of cluster_1. While this theoretical observation of 3D-connected nanoparticles still demands experimental verification, we highlight that it seems likely considering the high density of nanoparticles observed in the experiment [39], which results from the chemical complexity in HEAs. We further argue that, if the nanoparticles are indeed connected, it can significantly impact the motion of dislocations, which then have to go through the ordered nanoparticles rather than circumvent them. Other than the mechanical properties, the morphology of the nanoparticles also have important application in catalysis, which would be an interesting topic for future research.

### v2

@@ 2.2 Energy model
Figure 2: The accuracy of the energy model and the calculated specific heats for FeCoNiAlTi and MoNbTaW. (a) The training error of the EPI model for FeCoNiAlTi. (b) The validation error of the EPI model for FeCoNiAlTi (0.7/0.3 train-valid splitting). (c) The training error of the loca_SRO model for FeCoNiAlTi. (d) The validation error of the local_SRO model for FeCoNiAlTi. (e) The specific heats $C_{V}$ of FeCoNiAlTi were calculated with different supercell sizes. (f) The specific heats of MoNbTaW were calculated with a one-million-atom supercell.
@@
In this work, we focus on two types of energy models: effective-pair-interaction (EPI) model [59, 48], and a nonlinear model that uses the local SRO parameter as the input feature. The EPI model is a generalized Ising model constructed via machine learning from the DFT data. By automatically selecting the interaction range via Bayesian information criterion (BIC) [48], the EPI model has demonstrated that it can predict the DFT-calculated configuration energies of a series of HEAs with very high accuracy [33]. The local-SRO model, proposed in this work, generalizes the EPI model by adding a quadratic term. As will be shown in the following discussion, this simple nonlinear term has a profound impact on the computing pattern, and serves as a prototypical representation for machine learning models. A detailed description of the two energy models are present in Method 3.3.
@@
Using a train-validation splitting of 70% and 30% for the DFT dataset described in Method. 3.4, we evaluated the accuracy of the second-nearest-neighbor (2nd-NN) EPI model and the loca_SRO model and the results are shown in Fig. 2 (a-d). Both models are trained with the Adam optimizer and back-propagation algorithm. It can be seen that both of the two models demonstrate very high accuracy. For the EPI model, the validation root-mean-square-error (RMSE) is 0.1819 mRy, or approximately 2.5 meV, which is well within quantum chemical accuracy (approximately 3.16 mRy/atom), as well as smaller than the typical errors of $5-12$ meV/atom in the DFT methods [60]. The local_SRO model demonstrates slightly higher accuracy than the EPI model, with a validation RMSE of 0.1643 mRy/atom, which can be attributed to its inclusion of high-order interactions. The high accuracy of the models is also demonstrated in the $R^{2}$ scores, which are all higher than 0.995. The errors from the train and validation datasets are also very close to each other, indicating the absence of overfitting in the models. For comparison, we notice that this accuracy is higher than that in Ref. [29], which is reported as 14.7 meV/atom by applying a similar pair-interaction model in the FeCoNiAlTiB system. We think one reason for the enhanced accuracy is our approach to obtaining the DFT dataset, which combines both random configurations, as well as configurations from Monte Carlo simulations, as proposed in Ref. [33] and described in Method. 3.4.
@@
Using the trained models, we calculated the specific heats $C_{V}$ in the FeCoNiAlTi HEA in temperatures ranging from 50 K to 2000 K. The calculation details are given in Method. 3.5, and the results are shown in Fig. 2 (e). It can be seen that there is a sharp order-disorder transition at a low temperature of about 300 K. The sharp peak evolves to a singularity with the increase of the system size, which signifies the occurrence of a first-order transition. We also calculated the $C_{V}$ curve for the MoNbTaW refractory HEA. The results in Fig. 2 (f) agree well with the results in Ref. [33], which are calculated using a small system of 1000 atoms. Note that the order-disorder transition at lower temperatures (between 250 K and 500 K) in the $C_{V}$ curve generally cannot be directly compared with experiments due to kinetic barrier effects. On the other hand, the results at elevated temperatures are more suitable for comparison if experimental data are available, as demonstrated in Ref. [34]. A comparison of Fig. 2 (e) and (f) in the temperature range between 500 K and 2000K shows an interesting difference between the two HEAs: The $C_{V}$ curve in $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ is generally flat but demonstrates clear fluctuation as the temperature changes, indicating the existence of complex second phases or short-range order. By comparison, the $C_{V}$ curve of MoNbTaW in the same temperature region is smooth, but with an obvious order-disorder transition around 1000 K. Note that the exact location of the order-disorder transition temperature depends on the energy model and different values, as having been extensively reported in various theoretical works [61, 28, 33, 62].
2.3 Nanostructures in $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$
@@
Figure 3: Simulation results for FeCoNiAlTi with one-million atoms. (a) A snapshot of the (001) face at 1000 K. (b) a perspective 3D snapshot of the configuration at 1000 K, from the signified direction (c) Slices of the atomic system to show the nanoparticle, using a step of 10 fcc layers. (d) A snapshot view of the (001) face at 2000 K (e) a 3D view of the configuration at 2000 K. (f) High-resolution atom maps showing the atomistic distribution within the $L1_{2}$ nanoparticles of the Al7Ti7 alloy. Reproduced from Fig. 2 in Ref. [39]. (g) TEM image of the Al7Ti7 alloy showing the nanostructured morphology. Reproduced from Fig. 1 in Ref. [39] with reprint permission from Science.
@@
In this section, we report the nanostructure evolution of the $\mathrm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ fcc HEA. This alloy has demonstrated exceptional combined strength and ductility in experiments, and such an attractive property has been attributed to the formation of nanoparticles in the FeCoNi matrix [39], which hinder the motion of dislocations. In the experiment, the alloy sample is homogenized at 1150 ${}^{\circ}C$ (1423 K) for 2 hours and aged at 780 ${}^{\circ}C$ (1053 K) for 4 hours [39]. Using the local SRO model, we first simulated the chemical phase changes of the $\mathrm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ fcc alloy using a supercell of approximately one million atoms. The lattice dimension is chosen as $72\times 60\times 60$, therefore, the total number of atoms is $72\times 60\times 60\times 4=1,036,800$. We initialize the configurations randomly at 2000 K, then decrease the simulation temperature by 50 K each time until 1000 K, which is approximately the temperature at which the samples are aged. At each temperature, we skip the $4\times 10^{5}-1.5\times 10^{6}$ MC sweeps before recording the following 20,000 sweeps to calculate the specific heat $C_{V}$ in Fig. 2 (e). The configurations at 2000 K and 1000 K are demonstrated in Fig. 3 (a-e). From Fig. 3 (d-e) it can be seen that the system is chemically disordered at 2000 K. As the temperature decreases to 1000 K, a nanoparticle (NP) forms in the system. The distribution of elements in the NP, as shown in Fig. 3 (a-b), reveals the formation of $L1_{2}$ structure. The two different layers of the $L1_{2}$ nanoparticle can be seen from Fig. 3 (b-c). This agrees well with the experimental results in Ref. [39], as reproduced in Fig. 3 (f-g), in which the nanoparticles are refered to as multi-component intermetallic nanoparticles (MCINP). Similar results are also reported in Ref. [63] for the $\rm{Fe_{25}Co_{25}Ni_{25}Al_{15}Ti_{10}}$ HEA, where the fcc phase consisted of a $\gamma$ Fe-(Co,Ni)-based solid-solution matrix (A1), and coherent primary $\gamma^{\prime}$ (Ni,Co)3-(Ti,Al)-based intermetallic $L1_{2}$ precipitates. Other than chemical structure, size is another important feature of the nanoparticles. The diameter of the cluster shown in Fig. 3 (a) is about 17.8 nm, which is in agreement with the experimental observation shown in Fig. 3 (f-g). Moreover, we extract the compositions from the simulation data and show it in Fig. 4 (g), along with the experiment results in Ref. [39], as shown in Fig. 4 (h). It can be seen that the compositions from MC simulation generally agrees well with experiment, although there are also some differences, such as slightly higher concentration of Ni in MCINP.
@@
Figure 4: Comparing the simulated configurations of the 1 billion atom FeCoNiAlTi with the experimental results. (a) A histogram for all clusters of sizes larger than 100 atoms. Five clusters of different sizes are selected and shown as vertical dashed lines. (b) The chemical concentrations of the matrix (cluster_0) and cluster_1, as well as the rest clusters. (c-f) Snapshots of the [001] face of NiCoFeAlTi at T=1000K, with (c) as the enlarged upper-right corner of (d), and (e) as the enlarged upper-right corner of (f). (c-d) show the different elements, and (e-f) show the atomic local energies. (g) Compositions of the NP (MCINP) and matrix (MCM) from the 1M atom simulation via sampling along the x direction using a radius of 6.4 nm. (h) Compositions of the NP (MCINP) and matrix (MCM) from experiment, as reproduced from Ref. [39].
Figure 5: Comparison between the NPs from MC simulation and APT experiment. (a-c) 3D snapshot of cluster_4, 3, and 2, respectively. (d) APT results from fcc grains of FNAT sample (47-hours aging at 973 K). (e) Same as (d), but for a sample obtained via 4-hours aging. Both figure (d) and (e) are removed and shown as blank on purpose for permission reasons in this preprint. See original ones at Fig. 2 of Ref. [40]. (f) The APT results of the $\mathrm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ HEA, as reproduced from [39]. (g) The simulation result obtained using SMC. The blue region show the NP identified, and the red region show the matrix phase. (h) The simulation result obtained with SMC. The chemical species are explicitly shown.
@@
From Fig. 3 (b,c), we see that the one-million-atom supercell can only accommodate one NP in the FeCoNiAlTi HEA, and larger systems are necessary for a direct comparison of experimental data that contains more than one NP, such as the TEM image in Fig. 3 (g). To investigate the microstructures of the FeCoNiAlTi HEA at a larger length scale, we employ a $630\times 630\times 630$ fcc supercell, in which the total number of atoms is 1,000,188,000. Due to the huge supercell size, we use the faster EPI model rather than the local SRO model, and the simulation details are specified in Method. 3.5. The simulation results for this one billion-atom system are presented in Fig. 4 and Fig. 5. To study the size, composition, and morphologies of the nanoparticles, we employ the union-find algorithm to identify the nanoparticles, as detailed in Method. 3.5. We refer to all sets of atoms identified by the union-find algorithm as clusters, which can be of sizes ranging from a single isolated atom to the whole matrix phase. We make a histogram of all clusters of sizes larger than 100 atoms in Fig. 4 (a). The largest one is cluster_0, which is the matrix phase. Cluster_0 has a total of 0.6 billion atoms, and contains only Fe, Co, and Ni, as demonstrated in Fig. 4 (b). Note that the absence of Al and Ti atoms in cluster_0 is a result of the union-find algorithm, which automatically identifies isolated Al and Ti atoms in cluster_0 as individual clusters of size one. From Fig. 4 (c), we see that the FeCoNi matrix phase is a disordered A1 structure. The cause of the disorder can be seen from Fig. 4 (e) and (f), which show the local energies of each atom with the color bar. It can be seen that in the matrix phase, the local energies of different atoms are generally close to each other, which means that the Fe, Co, and Ni atoms have no significant site preferences. On the other hand, the nanoparticles, which forms $L1_{2}$ structure as shown in Fig. 4 (c), contains sites with much lower atomic local energies, as shown in Fig. 4 (e). In fact, the local energy difference is exactly the criteria we make use of to distinguish the matrix and nanoparticles, as described in Method 3.5. The FeCoNi matrix phase is the entropy-stablized disordered phase, while the $L1_{2}$ phase is the enthalpy favorable ordered phase. In other words, the occurrence of both disordered A1 structure and ordered $L1_{2}$ structure is a result of the competition between the entropy and enthalpy contributions in the free energy.
@@
To study the size, compositions, and morphologies of the nanoparticle in more detail, we choose four representative nanoparticles, as illustrated in Fig. 4 (a) as cluster_1, 2, 3, and 4, which have 264,792,042, 1,007,916, 102,389, and 20,937 atoms. Their chemical concentrations are also listed in Tab. 2, which aligns well with the experimental results [29]. From Fig. 4 (a), we note that the peak of the histogram is at about $10^{5}$ atoms, which corresponds to a cluster of 29 atoms, or 10.3 nm in diameter. For cluster_1, a similar calculation gives a value of 22.0 nm in diameter, if we assume that the particles are spherical. However, inspecting the 3D morphology of the clusters reveals that they are actually not simply spheres or ellipsoids, as shown in Fig. 5 (a-c): While ellipsoid is a good approximation for the smaller cluster_4, the larger cluster_3 and cluster_2 are both comprising of multiple nanoparticles, which indicates that the smaller nanoparticles can merge together to form larger ones, as shown by the dendritic structures at the surface of the nanoparticles. In fact, from Fig. 4 (a,b) we see that the largest $L1_{2}$ cluster: cluster_1, contains a total of more than 264 million atoms, which accounts for 66% of the atoms in the NP phases. In other words, most of the nanoparticles shown in Fig. 4 (c-f) are actually connected at 3-dimension and are parts of cluster_1. Compared to simple NP, we propose that this 3D-connected-NPs (3DCNP) can further hinder the motion of dislocations, which can serve as a strengthening mechanism in these types of HEAs. Verifying the existence of the 3DCNP demands high-precision experimental techniques such as atom probe tomography (APT). The APT image of $\rm{Fe_{29}Co_{29}Ni_{28}Al_{7}Ti_{7}}$ is available in Ref. [39], as reproduced in Fig. 5 (f). However, the image only shows the surface of the sample instead of the 3D chemical structure. Although the theoretical observation of the 3DCNP still requires experimental validation, we contend that its existence is plausible, given the high density of NPs observed in experiment [39]. Moreover, the APT result of the FNAT alloy [40], a medium entropy alloy made up of the Fe, Ni, Al, and Ti elements, seems to support the existence of $L1_{2}$ 3DCNP, as shown in Fig. 5 (d,e). For a direct comparison with the experimental results, we also show the simulation results of the APT sample needle in Fig. 5 (g, h). It is easy to see that the general shapes of the NPs from simulation resemble the experimental results of FNAT, despite their different chemical compositions. The sizes of the NPs from simulation is generally smaller than experimental results, which could be due to the relatively limited $10^{5}$ MC simulation steps. Again, the importance of SMC method is highlighted by the capability to directly simulate an APT needle comprising of 10 million atoms. Other than the mechanical properties, the NPs can also have important application in catalysis [64], which would be an interesting topic for future research [46, 65].